OpenStudy (anonymous):

Solve the equation. If it is an identity or contradiction, so state. Please show all work. x-5+6x=-2+7x

jimthompson5910 (jim_thompson5910):

x-5+6x=-2+7x 7x-5=-2+7x 7x=-2+7x+5 7x-7x=-2+5 0x=-2+5 0x=3 0=3 Since this equation is NEVER true for any x value, this means that there are no solutions.

OpenStudy (anonymous):

contradiction